OpenStudy (anonymous):

Find two consecutive even integers such that the smaller added to two times the larger gives a sum of 40.

OpenStudy (yttrium):

x+2(x+2) = 40 x+2x+4 = 40 3x = 36 x = 12 --> ans x+2 = 14 --> ans Chechking: 12+2(14) = 12+28 = 40 (correct)

OpenStudy (anonymous):

thank for the help